# SAS: Simulated Attention Score

### v1

@@ 1 Introduction
Building on these observations and insights, we propose Simulated Attention Score (SAS) to enhance Transformer performance through simulated attention head and feature expansion. Traditional query embeddings have dimensionality $[B,T,H,D]$, where $B$ represents the batch size, $T$ is the sequence length, $H$ denotes the head count, and $D$ is the feature dimension per head. Our approach applies linear projections with nonlinear activations along the head dimension ($H$) to effectively increase the number of attention heads to $\hat{H}$ with $\hat{H}>H$. Similarly, we also simulate more feature dimensions ($D$) for query/key embedding for the attention score computation. The proposed method effectively and efficiently expands both the head and feature dimensions, resulting in a query representation of shape $[B,T,\hat{H},\hat{D}]$, where the expanded dimensions $\hat{H}$ and $\hat{D}$ exceed the original dimensions $H$ and $D$, respectively. This expansion simulates a greater number of attention heads and a higher hidden feature dimensionality per head, thereby enhancing model expressiveness while preserving a compact overall model size. Additionally, we propose the parameter-efficient attention aggregation to control parameter size and computation. To summarize, our main contributions are the following:
@@ 2 Related Works
The Transformer architecture, leveraging Multi-Head Attention (MHA) (vaswani2017attention, ) for effective long-range dependency modeling, suffers from high memory footprint and bandwidth demands during training. To alleviate these issues, several efficient attention mechanisms have been proposed. Multi-Query Attention (MQA) (shazeer2019fast, ) reduces memory usage by sharing key and value heads among multiple queries. Grouped-Query Attention (GQA) (ainslie2023gqa, ) addresses MQA’s drawbacks by grouping queries with shared key and value heads. Multi-Head Latent Attention (MLA) (liu2024deepseek, ) optimizes inference through low-rank key and value representations, surpassing MHA with a smaller KV cache. Multi-matrix Factorization Attention (MFA) (hu2024multi, ) builds upon MQA by factorizing the query matrix. Tensor Product Attention (TPA) (zhang2025tensor, ) achieves performance improvements and KV cache reduction by employing contextual tensor decomposition for query, key, and value representations.
@@
Linear recurrent language models have emerged as a compelling alternative due to their training and inference efficiency. The evolution of these models has progressed from early linear RNNs with data-independent decay, exemplified by S4 (guefficiently, ), S5 (smithsimplified, ), LRU (orvieto2023resurrecting, ), RWKV4/5 (peng2023rwkv, ; peng2024eagle, ), and RetNet (sun2023retentive, ), towards more sophisticated architectures incorporating data-dependent decay, such as HGRN (qin2023hierarchically, ), Mamba (gumamba, ; daotransformers, ), and GSA (zhang2024gated, ). This transition highlights the effectiveness of gating and forgetting mechanisms (termed “selective mechanisms” in Mamba), a principle inspired by classical gated RNNs. A key distinction in modern forget gates, unlike those in LSTMs (graves2012long, ), is their independence from the previous hidden state, relying solely on input data and enabling efficient parallelization (qin2023hierarchically, ). Further advancements include Longhorn (liu2024longhorn, ) and TTT (sun2024learning, ), which reformulate state space learning as a gradient-based online optimization, and Gated DeltaNet (yang2024gated, ), which enhances a linear RNN’s expressiveness through gating while preserving training efficiency. sieberdifferenceFoundModels provides a comprehensive comparison of expressivity between Attention-based architectures, Linear attention mechanisms, and Recurrent Neural Networks.
@@ 3.2 Simulated Attention Score
Inspired by the above observations, we propose Simulated Attention Score (SAS), a method that simulates the behavior of a larger Transformer model with more attention heads and larger hidden size per head, using a smaller and more efficient model. Our approach maintains a moderate model size while significantly enhancing representational expressiveness. The implementation is presented in Appendix I.
@@
Theoretically, our method is compatible with various existing attention mechanisms, including MQA, GQA, MHA, and TPA (e.g. the result presented in Appendix G), to construct the initial query and key tensors. Among these, MHA offers relatively stable performance, and it is widely used in previous works vaswani2017attention ; kaplan2020scaling . Therefore, we adopt MHA as the foundation for initializing the query and key embeddings in our design.
@@ 3.3 Simulating Attention Heads via Non-linear Mappings
where $\text{Linear}_{1}^{Q_{h}}:\mathbb{R}^{H}\to\mathbb{R}^{\hat{H}}$ and $\text{Linear}_{1}^{Q_{h}}:\mathbb{R}^{\hat{H}}\to\mathbb{R}^{\hat{H}}$ are general linear transformation layers. Here, $H$ denotes the original number of attention heads, and $\hat{H}$ represents the expanded (simulated) head dimension. To improve training stability and mitigate the risk of vanishing gradients introduced by the transformation, we incorporate a residual (skip) connection from $\mathbf{Q}_{1}$ to $\hat{\mathbf{Q}}$. The additional parameter cost introduced by this transformation is negligible, as both the original head dimension $H$ and the expanded head dimension $\hat{H}$ are significantly smaller than the feature dimension $D$. Consequently, in auto-regressive Transformer models, the majority of the parameters are concentrated in the processing of feature vectors rather than in the introduced transformations along head dimensions.
@@ 3.5 Parameter-Efficient Attention Aggregation (PEAA)
for $i=1,2,\cdots,\hat{H}/H$, where we set $\hat{H}/H$ as a positive integer, and $\mathbf{W}^{O}\in\mathbb{R}^{(H\cdot D)\times(H\cdot D)}$ is responsible for a linear projection. In general, the attention feature aggregation is conducted by concatenating the output $\mathbf{h}_{i}$ among all expanded heads followed by a linear projection. However, due to the expanded (simulated) heads, it will introduce additional parameters and computation overhead. Here, we adopt the PEAA method that aggregates the attention features by averaging over groups of heads, as shown in Equation 12.
@@ 4.4 The Performance on Large-Scale Pretrain Dataset
Downstream Evaluation. We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (clark2018think, ), HellaSwag (zellers2019hellaswag, ), PIQA (bisk2020piqa, ), ScIQ welbl2017crowdsourcing , SocialIQA sap2019social and WinoGrande (sakaguchi2021winogrande, ), using the lm-evaluation-harness codebase (eval-harness, ). For ARC-E, ARC-C, HellaSwag, PIQA, and ScIQ, we report accuracy norm; for other tasks, we report standard accuracy. We display the zero-shot evaluation results of models here in Tables 1.
@@ Appendix C Model Configuration
Table 2: The model configuration of different models follows the setting of TPA [93].
@@
MHA 768 16 64 -
MQA 768 31 64 31
GQA 768 30 64 15
MLA 768 23 64 -
TPA 768 48 64 -
SAS 768 48 96 -
For the experiment on the Fineweb-Edu dataset, the experiment setup is as follows: We follow the nanoGPT training configuration. In particular, we use the AdamW [44] optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$. We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler [43]. For the model setting, we mostly follow the setting of TPA [93].
@@ Appendix F The Training Cost
Table 4: The Time cost and Memory cost of different methods with training length 512 and batch size 1.
Model 125M 125M 350M 350M 2.7B 2.7B
Time (ms) Memory (GB) Time (ms) Memory (GB) Time (ms) Memory (GB)
MHA 107.05 1.52 216.13 3.69 544.23 15.97
GQA 110.89 1.57 225.91 3.88 706.57 17.46
MQA 111.91 1.57 227.55 3.87 708.93 17.43
MLA 180.73 1.53 331.29 4.55 1071.56 20.17
TPA 125.04 1.53 256.23 3.86 904.71 17.77
SAS 205.76 1.53 419.35 3.72 736.73 16.14
We present a comprehensive analysis of the time and memory costs for various attention mechanisms across varying model sizes, as shown in Table 6. Here, the training time reported corresponds to the seconds for each training step. Across all model sizes, the MHA demonstrates competitive time performance, serving as a strong baseline. For the 2.7B parameter model, MHA completes training in 544.23ms. For the larger model (e.g., 2.7B model ), compared to MLA and TPA (recently proposed methods), SAS shows better time efficiency while maintaining competitive memory usage. Relative to the MHA baseline, SAS requires only minimal additional resources, suggesting its practical viability as a drop-in replacement. The advantage of SAS becomes more pronounced at larger model scales, demonstrating better scalability properties than several competing approaches. These findings suggest that SAS offers a compelling trade-off between computational efficiency and memory consumption, particularly for large-scale models where both time and memory resources are critical constraints.
Appendix G SAS with Different Attention Types
@@
Appendix H The Experiment Result on Large Model and Dataset
Table 6: The performance of different models with 1.5B model and 50B tokens. The performance of MHA, MQA, GQA, MLA and TPA comes from TPA paper [93]
@@ Appendix H The Experiment Result on Large Model and Dataset
Appendix I Implementation
In this section, we present the implementation of the proposed SAS module in PyTorch for research purposes, which is consistent with the intended use [49].
@@ Appendix I Implementation
class ResidualCNN(nn.Module):
@@
self.ReLU = nn.ReLU()
self.process = nn.Conv1d(input, output, kernel_size, 1, padding)
@@
class ResidualMLP(nn.Module):
@@
self.ReLU = nn.ReLU()
self.process = nn.Linear(input, output)
@@
self.sas_q = nn.Sequential(
nn.Conv1d(self.ori_head, self.target_head, kernel_size, 1, padding),
@@
nn.Linear(self.ori_feature, self.target_feature), ResidualMLP(self.target_feature, self.target_feature))
self.sas_k = nn.Sequential(
nn.Conv1d(self.ori_head, self.target_head, kernel_size, 1, padding),
@@
nn.Linear(self.ori_feature, self.target_feature), ResidualMLP(self.target_feature, self.target_feature))
self.sas_v = nn.Sequential(
nn.Conv1d(self.ori_head, self.target_head, kernel_size, 1, padding),
@@
self.output_dense=nn.Linear(self.ori_head*self.ori_feature,self.ori_head*self.ori_feature)

### v2

@@ 1 Introduction
Building on these observations and insights, we propose Simulated Attention Score (SAS) to enhance Transformer performance through simulated attention head and feature expansion. Traditional query embeddings have dimensionality $[B,T,H,D]$, where $B$ represents the batch size, $T$ is the sequence length, $H$ denotes the head count, and $D$ is the feature dimension per head. Our approach applies linear projections with nonlinear activations along the head dimension ($H$) to effectively increase the number of attention heads to $\hat{H}$ with $\hat{H}>H$. Similarly, we also simulate more feature dimensions ($D$) for query/key embedding for the attention score computation. The proposed method effectively and efficiently expands both the head and feature dimensions, resulting in a query representation of shape $B,T,\hat{H},\hat{D}$], where the expanded dimensions $\hat{H}$ and $\hat{D}$ exceed the original dimensions $H$ and $D$, respectively. This expansion simulates a greater number of attention heads and a higher hidden feature dimensionality per head, thereby enhancing model expressiveness while preserving a compact overall model size. Additionally, we propose the parameter-efficient attention aggregation to control parameter size and computation. To summarize, our main contributions are the following:
@@ 2 Related Works
The Transformer architecture, leveraging Multi-Head Attention (MHA) (vaswani2017attention) for effective long-range dependency modeling (xiong2024uncomp; xiong2025parallelcomp), suffers from high memory footprint and bandwidth demands during training. To alleviate these issues, several efficient attention mechanisms have been proposed. Multi-Query Attention (MQA) (shazeer2019fast) reduces memory usage by sharing key and value heads among multiple queries. Grouped-Query Attention (GQA) (ainslie2023gqa) addresses MQA’s drawbacks by grouping queries with shared key and value heads. Multi-Head Latent Attention (MLA) (liu2024deepseek) optimizes inference through low-rank key and value representations, surpassing MHA with a smaller KV cache. Multi-matrix Factorization Attention (MFA) (hu2024multi) builds upon MQA by factorizing the query matrix. Tensor Product Attention (TPA) (zhang2025tensor) achieves performance improvements and KV cache reduction by employing contextual tensor decomposition for query, key, and value representations.
@@
Linear recurrent language models have emerged as a compelling alternative due to their training and inference efficiency. The evolution of these models has progressed from early linear RNNs with data-independent decay, exemplified by S4 (guefficiently), S5 (smithsimplified), LRU (orvieto2023resurrecting), RWKV4/5 (peng2023rwkv; peng2024eagle), and RetNet (sun2023retentive), towards more sophisticated architectures incorporating data-dependent decay, such as HGRN (qin2023hierarchically), Mamba (gumamba; daotransformers), and GSA (zhang2024gated). This transition highlights the effectiveness of gating and forgetting mechanisms (termed “selective mechanisms” in Mamba), a principle inspired by classical gated RNNs. A key distinction in modern forget gates, unlike those in LSTMs (graves2012long), is their independence from the previous hidden state, relying solely on input data and enabling efficient parallelization (qin2023hierarchically). Further advancements include Longhorn (liu2024longhorn) and TTT (sun2024learning), which reformulate state space learning as a gradient-based online optimization, and Gated DeltaNet (yang2024gated), which enhances a linear RNN’s expressiveness through gating while preserving training efficiency. sieberdifferenceFoundModels provides a comprehensive comparison of expressivity between Attention-based architectures, Linear attention mechanisms, and Recurrent Neural Networks.
@@ 3.2 Simulated Attention Score
Inspired by the above observations, we propose Simulated Attention Score (SAS), a method that simulates the behavior of a larger Transformer model with more attention heads and larger hidden size per head, using a smaller and more efficient model. Our approach maintains a moderate model size while significantly enhancing representational expressiveness. The implementation is presented in Appendix L.
@@
Theoretically, our method is compatible with various existing attention mechanisms, including MQA, GQA, MHA, and TPA (e.g. the result presented in Appendix J), to construct the initial query and key tensors. Among these, MHA offers relatively stable performance, and it is widely used in previous works vaswani2017attention; kaplan2020scaling. Therefore, we adopt MHA as the foundation for initializing the query and key embeddings in our design.
@@ 3.3 Simulating Attention Heads via Non-linear Mappings
where $\text{Linear}_{1}^{Q_{h}}:\mathbb{R}^{H}\to\mathbb{R}^{\hat{H}}$ and $\text{Linear}_{2}^{Q_{h}}:\mathbb{R}^{\hat{H}}\to\mathbb{R}^{\hat{H}}$ are general linear transformation layers. Here, $H$ denotes the original number of attention heads, and $\hat{H}$ represents the expanded (simulated) head dimension. To improve training stability and mitigate the risk of vanishing gradients introduced by the transformation, we incorporate a residual (skip) connection from $\mathbf{Q}_{1}$ to $\hat{\mathbf{Q}}$. The additional parameter cost introduced by this transformation is negligible, as both the original head dimension $H$ and the expanded head dimension $\hat{H}$ are significantly smaller than the feature dimension $D$. Consequently, in auto-regressive Transformer models, the majority of the parameters are concentrated in the processing of feature vectors rather than in the introduced transformations along head dimensions.
@@ 3.5 Parameter-Efficient Attention Aggregation (PEAA)
for $i=1,2,\cdots,\hat{H}/H$, where we set $\hat{H}/H$ as a positive integer, and $\mathbf{W}^{O}\in\mathbb{R}^{(H\cdot D)\times(H\cdot D)}$ is responsible for a linear projection. In general, the attention feature aggregation is conducted by concatenating the output $\mathbf{h}_{i}$ among all expanded heads followed by a linear projection. However, due to the expanded (simulated) heads, it will introduce additional parameters and computation overhead. Here, we adopt the PEAA method that aggregates the attention features by averaging over groups of heads, as shown in Equation˜12.
@@ 4.4 The Performance on Large-Scale Pretrain Dataset
Downstream Evaluation. We evaluate zero-shot and two-shot performance on standard benchmarks, including ARC (clark2018think), HellaSwag (zellers2019hellaswag), PIQA (bisk2020piqa), ScIQ welbl2017crowdsourcing, SocialIQA sap2019social and WinoGrande (sakaguchi2021winogrande), using the lm-evaluation-harness codebase (eval-harness). For ARC-E, ARC-C, HellaSwag, PIQA, and ScIQ, we report accuracy norm; for other tasks, we report standard accuracy. We display the zero-shot evaluation results of models here in Tables 1.
@@ Appendix C Model Configuration
Table 2: The model configuration of different models follows the setting of TPA [zhang2025tensor].
@@
MHA 1024 16 64 -
MQA 1024 31 64 31
GQA 1024 30 64 15
MLA 1024 23 64 -
TPA 1024 48 64 -
SAS 1024 48 96 -
For the experiment on the Fineweb-Edu dataset, the experiment setup is as follows: We follow the nanoGPT training configuration. In particular, we use the AdamW [loshchilov2019decoupled] optimizer with $(\beta_{1},\beta_{2})=(0.9,0.95)$, a weight decay of $0.1$, and gradient clipping at $1.0$. We follow the same setting as nanoGPT that the learning rate is managed by a cosine annealing scheduler [loshchilov2016sgdr]. For the model setting, we mostly follow the setting of TPA [zhang2025tensor].
@@ Appendix F The Training Cost
Table 4: The Time cost of different methods with training length 512 and batch size 1.
Model 125M 350M 2.7B 6.7B 10.6B
MHA 36.20 72.87 117.82 186.58 217.01
GQA 38.87 71.66 133.76 205.69 239.16
MQA 39.37 72.01 132.62 202.61 238.85
MLA 51.68 104.95 180.03 254.43 637.78(Zero3 to avoid OOM;H200)
TPA 44.19 86.56 151.71 235.85 258.41
SAS 68.07 138.53 215.13 350.70 402.15
Appendix G Large Mode Size 6.7B and 10.6B performance
Table 5: The perplexity of 6.7B and 10B on the Pile dataset.
Model 5K 10K 15K 20K 25K 30K 35K 40K 45K 50K
6.7B
MHA 23.45 17.25 14.51 13.21 12.07 11.21 10.52 10.28 9.91 9.73
GQA 21.62 16.35 13.93 12.74 11.69 10.88 10.24 10.01 9.65 9.49
MQA 21.56 16.32 13.93 12.73 11.70 10.90 10.24 10.02 9.66 9.49
MLA 22.22 16.67 14.26 13.05 11.99 11.19 10.53 10.29 9.91 9.74
SAS 20.44 15.32 13.17 12.15 11.21 10.51 9.93 9.73 9.40 9.25
10.6B
MHA 23.67 17.34 14.59 13.18 12.00 11.15 10.46 10.20 9.82 9.64
GQA 21.86 16.45 13.92 12.70 11.61 10.81 10.17 9.94 9.58 9.41
MQA 22.13 16.55 14.03 12.80 11.70 10.88 10.21 9.98 9.61 9.43
MLA 22.40 16.67 14.22 12.99 11.90 11.11 10.42 10.18 9.83 9.66
SAS 20.53 15.12 12.92 11.90 10.98 10.29 9.71 9.51 9.18 9.04
Appendix H Large Training Length performance
Table 6: The perplexity of length 16384 on the Pile dataset
Model 5K 10K 15K 20K 25K 30K 35K 40K 45K 50K
MHA 18.31 14.50 12.85 11.96 11.45 11.11 10.44 10.37 9.95 9.94
GQA 17.74 14.40 12.86 11.96 11.47 11.14 10.45 10.39 9.99 9.98
MQA 18.02 14.46 12.82 11.96 11.44 11.10 10.46 10.38 9.96 9.99
MLA 17.74 14.40 12.86 11.96 11.47 11.14 10.45 10.39 9.99 9.98
TPA 18.49 14.54 12.96 12.08 11.59 11.26 10.62 10.58 10.16 10.18
SAS 14.93 12.45 11.23 10.56 10.16 9.90 9.32 9.28 8.95 8.97
Appendix I Theoretical Analysis
The Softmax Attention with Fully Parameterized Bilinear Attention $O_{i}=\sum_{c=1}^{H}\left(\sum_{j=1}^{i}\phi(\frac{x_{i}W_{c}x_{j}}{\sqrt{D}})x_{j}U_{c}\right)$ where: $W_{c}\in\mathbb{R}^{D\times D}$ = learnable bilinear transformation for head $c$, $\phi$ = the kernel function such as softmax, $U_{c}\in\mathbb{R}^{D\times D}$ = value projection matrix, $H$ = number of attention heads, $D$ = hidden dimension per head. Each head computes a weighted sum of values, with weights determined by $x_{i}W_{c}x_{j}^{T}$. For $m$-dimensional $x_{i}$ and $n$-dimensional $x_{j}$, one potential relation pattern is related to $x_{i}^{m}x_{j}^{n}$. Therefore, at most $D^{2}$ relations or interactions).
The bilinear term $x_{i}W_{c}x_{j}^{T}=\sum_{m=1}^{D}\sum_{n=1}^{D}x_{i}^{(m)}W_{c}^{(m,n)}x_{j}^{(n)}$ can be expressed as $(x_{i}\otimes x_{j})\cdot\text{vec}(W_{c})$, where $\text{vec}(W_{c})\in\mathbb{R}^{D^{2}}$. With $H$ heads, the model can represent up to $D^{2}$ distinct relations [6]. We hope that the output of each head is a single relation but not mixed relations, so that we can utilize the relations independently (e.g., $\phi$ processes the relation independently). Assume $Rank(W_{c})=1$, and there is only one non-zero value in $W_{c}$. And there are $H=D^{2}=4$. $x_{i}=[a~b]$ $x_{j}=[c~d]$. $W_{1}=[1~0]~[0~0]$ $W_{2}=[0~1]~[0~0]$ $W_{3}=[0~0]~[1~0]$ $W_{4}=[0~0]~[0~1]$. The first head $W_{1}$ will gives the result of relation $ac$, the $W_{2}$, $W_{3}$, $W_{4}$ will give the result of $ad$, bcbc, bdbd. Therefore, this ($ac$, adad, bcbc, bdbd) suggests different attention patterns and relations. And they could be adjusted independently and not mixed (e.g., $\phi$ processes the relation independently), compared to only one head with $W_{c}=[1~1$ 11]1~1] that outputs the mix of $ac+ad+bc+bd$. With a large $D$, we can have more combinations. With a large $H$, we can capture more independent relations. If we increase $D$ (e.g., hidden size), we increase the maximum number of relations. If we increase the $H$, we increase the number of independent relations that are actually captured.
Appendix J SAS with Different Attention Types
@@ Appendix I Theoretical Analysis
Appendix K The Experiment Result on Large Model and Dataset
Table 8: The performance of different models with 1.5B model and 50B tokens. The performance of MHA, MQA, GQA, MLA and TPA comes from TPA paper [zhang2025tensor]
@@
Appendix L Implementation
In this section, we present the implementation of the proposed SAS module in PyTorch for research purposes, which is consistent with the intended use [paszke2019pytorch].
@@
class ResidualCNN($n$.Module):
@@
self.ReLU = $n$.ReLU()
self.process = $n$.Conv1d(input, output, kernel_size, 1, padding)
@@
class ResidualMLP($n$.Module):
@@
self.ReLU = $n$.ReLU()
self.process = $n$.Linear(input, output)
@@
self.sas_q = $n$.Sequential(
$n$.Conv1d(self.ori_head, self.target_head, kernel_size, 1, padding),
@@
$n$.Linear(self.ori_feature, self.target_feature), ResidualMLP(self.target_feature, self.target_feature))
self.sas_k = $n$.Sequential(
$n$.Conv1d(self.ori_head, self.target_head, kernel_size, 1, padding),
@@
$n$.Linear(self.ori_feature, self.target_feature), ResidualMLP(self.target_feature, self.target_feature))
self.sas_v = $n$.Sequential(
$n$.Conv1d(self.ori_head, self.target_head, kernel_size, 1, padding),
@@
self.output_dense=$n$.Linear(self.ori_head*self.ori_feature,self.ori_head*self.ori_feature)